OpenStudy (anonymous):

Find the volume of a sphere with a diameter of 4 feet. (Give your answer correct to 2 decimal places.)

OpenStudy (anonymous):

\[Volume=(4/3)*\pi*R ^{3}\], where 2R=diameter

OpenStudy (anonymous):

Volume of sphere = (4/3)*pi*r^3 r =d/2 r=4/2=2 V = (4/3)*pi*(2)^3

OpenStudy (anonymous):

is the answer 10.66

OpenStudy (anonymous):

my answer differs. You should check wether you need to get your volume in feet or in meters and convert before calculations if needed!

OpenStudy (anonymous):

its Feet